The differential equation dy/dx = 25+20x+40y+32xy



has an implicit general solution of the form F(x,y)=K



In fact, because the differential equation is separable, we can define the solution curve implicitly by a function in the form of F(x,y)=G(x)+H(y)=K



Find such a solution and then give the related functions requested for


F(x,y)=G(x)+H(y)=

1
Expert's answer
2022-03-11T03:32:32-0500

The given equation can be rewritten as, "\dfrac{dy}{dx} = (4x+5)(8y+5)".


The equation can be separated as, "\dfrac{dy}{8y+5} = (4x+5)dx"


Integrating, we get


"\dfrac{1}{8}\log (8y+5) = 2x^2 + 5x + c"


Therefore,

"F(x,y)=G(x)+H(y)= 2x^2 + 5x - \dfrac{\log (8y+5)}{8}".